If Individual II-2 has two children with an unaffected partner, what is the probability that they both will have the condition?Assume that the pedigree below illustrates an autosomal dominant condition. SS
Answer #1

Answer:

Assume that the pedigree below illustrates an autosomal dominant condition. aa 15

II-2 (Aa) x (aa) unaffected parent ----Couple

a
A Aa (affected)-1/2
a aa (unaffected)--1/2

The probability that they both will have the condition = 1/2 * 1/2 = 1/4